Operations Report – May 3rd

Crew 261 Operations Report 03-05-2023
SOL: 3
Name of person filing report: Julien Villa-Massone
Non-nominal systems:
· Toilet
· Remote observatory
Notes on non-nominal systems:
Toilet started to smell a lot (making some crew mates uncomfortable when on the lower deck). Mission Support provided guidance:
1) Let’s fill the remaining toilet tank with water and empty it.
2) We need to make sure there are no leaks in the toilet/toilet tank:
Please inspect area in the toilet room for any leaks
In the EVA prep room please remove the toilet tank cover and inspect underneath with the flashlight to see if there any leaks
Crew Engineer applied this guidance. No evidence of leak was found. Over the course of the day, smell reduced. The smell level in the lower deck is slightly noticeable. In the toilet, the smell still exists, but it is not unbearable like it was before emptying.
Following up, Mission Support informed us that a septic tank cleaning was ordered and they should be at the bast tomorrow. Air freshener was offered if if smell continues.
We appreciate the help of Mission Support on this matter.
Spirit rover used: No
Hours: (before EVA) 220.3
Beginning charge: (Before EVA) 100
Ending charge: (On return from EVA, before recharging) N/A
Currently charging: Yes
Opportunity rover used: Yes
Hours: 123.0
Beginning charge: 100
Ending charge: 78
Currently charging: Yes
Curiosity rover used: Yes
Hours: 232.9
Beginning charge: 100
Ending charge: 99
Currently charging: Yes
Perseverance rover used: No
Hours: 262.8
Beginning charge: 100
Ending charge: N/A
Currently charging: Yes
General notes on rovers: Nominal operation today
Summary of Hab operations: Nominal day (toilet issue described above)
WATER USE: rate of 30 gallons / sol. Trend of 28 Gallons/sol since mission start. We are now on target to finish the mission 2 gallons short. (So we are essentially on target). This means, if we continue with this trend, the tank will be at 118 gallons when the mission ends.
Water (static tank): 388 gallons remaining
Static tank pipe heater (on or off): off
Static tank heater (On or off): off
Toilet tank emptied: yes

Summary of internet: Nominal
Summary of suits and radios: Suits: nominal. Radios: nominal. For information, as mentioned by Mission Support during initial training, we noticed that Radio #2 was unable to transmit reliably to the repeater on channel 1 (at 1.5km distance south, at the ridge), so the EVA crew decided to use channel 2 locally, and the crew member with the other radio would use channel 1 to communicate to Hab. Habcomm crew decided to have one radio on channel 1 and one radio on channel 2 to increase situational awareness.
Summary of GreenHab operations: GreenHab door closed earlier due to the wind
WATER USE: 6 gal at 06:30 + 4 gal
Heater: Off
Supplemental light: Off
Harvest: None today
Summary of ScienceDome operations: Nominal day with mainly work by Crew Biologist. Biostimulation solution prepared. Added culture + media to Bioreactor – now operational. Set up microscope to show crew Spirulina tomorrow.
Dual split: Off
POWER report: Nominal operations today. Crew Engineer drew a diagram of the power system components to gain a better understanding. The chart assumes 3 connections due to some wiring being hidden and inaccessible, but all other connections are visible and have been recorded and charted on the diagram. Logs analyzed to a limited extent, producing this graph. Crew Engineer started modeling the battery SoC based on VDC. Next, will attempt modeling according to Watt-hours. Battery charging / discharging seems nominal – generator has not required since mission start.

Power System SoC comes from the power system and seems to be inaccurate, as the SoC remains low and decreasing while the battery has positive watts (is charging). This is why Crew Engineer is attempting to model SoC.
Summary of RAM operations: Crew Robotics Engineer continued assembly and adjustments of the robot following a first EVA near the base for initial testing. Crew Engineer cut a spare PVC pipe (using a hand saw) to serve as support for a failed step in the Hab staircase.
Summary of any observatory issues: Nothing to report
Summary of health and safety issues: Nominal day
Questions, concerns and requests to Mission Support: Grateful for the continued support received from Mission Support. Thanks extended from the entire crew.

Copyright © The Mars Society. All rights reserved. | Main Site